# Fodor and Pylyshyn’s Legacy – Still No Human-like Systematic Compositionality in Neural Networks

### v1

@@ Abstract
Strong meta-learning capabilities for systematic compositionality are emerging as an important skill for navigating the complex and changing tasks of today’s world. However, in presenting models for robust adaptation to novel environments, it is important to refrain from making unsupported claims about the performance of meta-learning systems that ultimately do not stand up to scrutiny. While Fodor and Pylyshyn famously posited that neural networks inherently lack this capacity as they are unable to model compositional representations or structure-sensitive operations, and thus are not a viable model of the human mind, Lake and Baroni recently presented meta-learning as a pathway to compositionality. In this position paper, we critically revisit this claim and highlight limitations in the proposed meta-learning framework for compositionality. Our analysis shows that modern neural meta-learning systems can only perform such tasks, if at all, under a very narrow and restricted definition of a meta-learning setup. We therefore claim that ‘Fodor and Pylyshyn’s legacy’ persists, and to date, there is no human-like systematic compositionality learned in neural networks.
@@ 1 Introduction
Thus, in this paper we argue that: Neural networks have not yet achieved learning systematic compositional abilities. Specifically, based on a case of effective criticism of Lake and Baroni’s framework, we outline how to argue, test, and train for systematic generalization and compositionality, and demonstrate the relevance of our position (c.f. Figure 1 for a schematic).
We develop this position as follows: (I) We identify Fodor and Pylyshyn’s main arguments in the context of the compositionality challenge for artificial neural networks and locate the nature of Lake and Baroni’s approach in refuting Fodor and Pylyshyn’s claims that neural networks cannot reliably develop compositional representations and structure-sensitive operations. (II) We show that within their setup, the model exhibits various non-systematic behaviors that can not be considered human-like and clearly violates structure-sensitivity. (III) We argue for several necessary aspects of training and evaluation of meta-learning systems to achieve and assess their systematicity, taking into account the relevance of compositional representations and structure-sensitive operations. (IV) We adapt the arguments of Fodor and Pylyshyn in light of the modern development of deep learning systems to argue for a future of models capable of learning symbolic representations for artificial cognition and representation learning.
@@ 3 Systematicity through Meta-Learning
GRAMMAR #133 (Lake and Baroni) | ||||||||||||||||||||||||||
DECODING (this paper; for readability) | ||||||||||||||||||||||||||
SUPPORT (Lake and Baroni) | ||||||||||||||||||||||||||
QUERY (Lake and Baroni) | ||||||||||||||||||||||||||
QUERY (Lake and Baroni) | ||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||
@@ 3.1 Examining the lack of compositionality
Lake and Baroni mention that generalization beyond training occurs only with respect to new combinations of three grammar rules from the same set of grammar rules used during training. However, if we consider the invariance under the atomic assignments of colors to language tokens and the mere labeling of operations, we find that 179/200 validation episodes have a combination of non-primitive grammar operations that were already present in the 100000 training episodes. (See Appendix A.1 for details.) Thus, even if the model achieves highly systematic results on the test episodes, this could be due to memorization of the experienced operation patterns and learning to extract the correct labels from the episode’s support examples. However, we can even show that there is non-systematic behavior within their repository of test episodes; we re-evaluate their pre-trained model on the same set of test episodes, with the only difference that we did 10 evaluations of all query examples for each test episode, for statistical purposes, similar to the one episode they further evaluated against human performance. We find that the model performs worst on episodes #133, #32, and #122, with accuracies of only 41%, 52%, and 54% on the query examples, respectively. (See the next paragraph and the Appendix A.2 for details.)
Failure in rule extraction. Further investigation of Episode #133 (see Table 1) reveals that the model has trouble correctly processing the semantics of the language token with the hidden grammar rule and will therefore call it and confuse it up with the token (with ) which we will call . It seems to have a problem with the only example with , , which also happens to contain . But since has several iconic examples in the support, it is expected that a reasoner with compositional skills will be able to systematically use a single example and remain consistent with the rest of the support information. Considering the examples , , , human systematicity would at least suspect some semantics of that are different from those of .
@@
GRAMMAR #1 (Lake and Baroni) |
DECODING (this paper; for readability) |
SUPPORT (Lake and Baroni) |
QUERY (new; this paper) | |||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||
Violating structure-sensitivity. Besides both previous failure modes that are related to incompetence in extracting information from the support examples, we also found query examples for episode #1 that reveal additional non-systematicity (see Table 2 or Appendix A.2 for extended version). For queries with the patterns and we first see that the model never parses as intended. Instead of and , the stable output can be explained by parsing as intended. Instead of and – except for the cases, , , , where it would make no difference! Only the (also ambiguous) case is processed correctly in 6/10 cases – but with even worse performance than with the unambiguous examples. Despite the structural similarity to the other query examples, down to the color combination, we see a non-systematic deviation in the response, which raises doubts about compositional skills.
@@ 5 Alternative Views
6 Position Summary and Discussion
For this final section, we will reiterate the key points that make up our position (see Sec. 1) and that we believe are important aspects of the goal of achieving meta-learning models capable of human-like systematic compositionality: (I) Criteria for compositionality. The main criteria for models with productive, systematic, and compositional capabilities remain compositional representation and structure-sensible operations. (II) Non-systematic behavior. Since Lake and Baroni’s model exhibits various non-systematic behaviors, it fails to demonstrate human-like compositional learning capabilities, and further refutes the presented claims that their meta-learning framework achieves human-like systematic generalization. (III) Assessment of compositionality. Systematic testing of multiple types of out-of-distribution episodes is necessary to assess compositional abilities and structure-sensitive operations. (IV) Emergence and Learning of Symbolic Representations. Meta-learning systems need to support the emergence of compositional symbolic representations during training. For this, we expect training tasks and model architectures that make iteration, self-validation, and self-correction over the extracted rule sets possible and necessary. The limitations of current neural models underscore the importance of hybrid architectures that integrate the strengths of symbolic and connectionist paradigms. Key advances in this direction include systematicity, reflective reasoning, and scalability.
@@ 7 Conclusion
While the importance of compositional representations and structure-sensitive operations for human-like systematicity remains, the previous consideration allows the training and testing of artificial neural networks that encourage the development of such properties. By bridging the gap between symbolic and connectionist principles, hybrid architectures may be particularly promising, since they do not suffer from the limitations of neural networks without symbolic machinery as specified by Fodor and Pylyshyn. Overall, the continued relevance of Fodor and Pylyshyn’s critique underscores the challenges of developing systems capable of systematic generalization and compositional reasoning. While meta-learning frameworks represent significant progress, they do not address fundamental limitations. Future advances must embrace integrative approaches that combine the strengths of symbolic and connectionist paradigms, paving the way for a more robust understanding of artificial cognition. By addressing these challenges, we can move closer to realizing the vision of human-like artificial intelligence.
@@
The authors would like to thank the reviewers of the ICML 2025 position paper track for their thoughtful feedback and the fruitful discussions. This work was funded by the European Union (Grant Agreement no. 101120763 - TANGO). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Health and Digital Executive Agency (HaDEA). Neither the European Union nor the granting authority can be held responsible for them. The authors acknowledge the support of the German Science Foundation (DFG) research grant "Tractable Neuro-Causal Models" (KE 1686/8-1) as well as the AlephAlpha Collaboration lab 1141. This work was also supported by the Priority Program (SPP) 2422 in the subproject “Optimization of active surface design of high-speed progressive tools using machine and deep learning algorithms“ funded by the German Research Foundation (DFG). Further, we acknowledge support of the hessian.AI Service Center (funded by the Federal Ministry of Education and Research, BMBF, grant No 01IS22091), the “Third Wave of AI”, and “The Adaptive Mind”.
@@ A.1 Note on analysis for different combinations of non-primitive grammar operations
In section 3.1 we state that only 179/200 validation episodes have a combination of non-primitive grammar operations that were already present in the 100000 training episodes. This is the result of counting every different combination of 3 operations, unary () or binary (), present in both training and validation episodes, when abstracting the individual function names .
@@ A.2 Extended outputs for Lake and Baroni’s meta-learning testing episodes
Below we include the full set of grammar rules, support examples, and query examples from Lake and Baroni’s meta-learning. We re-evaluated their pre-trained model on the same set of test episodes. Here we report the results for #133, #132, #122, and the modified #1.
@@
GRAMMAR #133 (Lake and Baroni) | ||||||||||||||||||||||||
DECODING (this paper; for readability) | ||||||||||||||||||||||||
SUPPORT (Lake and Baroni) | ||||||||||||||||||||||||
QUERY (Lake and Baroni) | ||||||||||||||||||||||||
QUERY (Lake and Baroni) | |||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||
@@
GRAMMAR #32 (Lake and Baroni) | |||||||||||||||||||||||||||||||||||
DECODING (this paper; for readability) | |||||||||||||||||||||||||||||||||||
SUPPORT (Lake and Baroni) | |||||||||||||||||||||||||||||||||||
QUERY (Lake and Baroni) | |||||||||||||||||||||||||||||||||||
QUERY (Lake and Baroni) | ||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||
@@
GRAMMAR #122 (Lake and Baroni) | |||||||||||||||||||||||||||||||||||||||||||
DECODING (this paper; for readability) | |||||||||||||||||||||||||||||||||||||||||||
SUPPORT (Lake and Baroni) | |||||||||||||||||||||||||||||||||||||||||||
QUERY (Lake and Baroni) | |||||||||||||||||||||||||||||||||||||||||||
QUERY (Lake and Baroni) | ||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||
@@
GRAMMAR #1 (Lake and Baroni) |
DECODING (this paper; for readability) |
SUPPORT (Lake and Baroni) |
QUERY (new; this paper) | |||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||

### v2

@@ Abstract
Strong meta-learning capabilities for systematic compositionality are emerging as an important skill for navigating the complex and changing tasks of today’s world. However, in presenting models for robust adaptation to novel environments, it is important to refrain from making unsupported claims about the performance of meta-learning systems that ultimately do not stand up to scrutiny. While Fodor and Pylyshyn famously posited that neural networks inherently lack this capacity as they are unable to model compositional representations or structure-sensitive operations, and thus are not a viable model of the human mind, Lake and Baroni recently presented meta-learning as a pathway to compositionality. In this paper, we critically revisit this claim and highlight limitations in the proposed meta-learning framework for compositionality. Our analysis shows that modern neural meta-learning systems can only perform such tasks, if at all, under a very narrow and restricted definition of a meta-learning setup. We therefore claim that ‘Fodor and Pylyshyn’s legacy’ persists, and to date, their main criteria for modeling human-like compositionality remain unchallenged or unsatisfied by modern approaches.
@@ 1 Introduction
Thus, in this paper we argue that: Neural networks have not yet achieved learning systematic compositional abilities. Specifically, based on a case of effective criticism of Lake and Baroni’s framework, we outline how to argue, test, and train for systematic generalization and compositionality (c.f. Figure 1 for a schematic).
We develop our argument as follows: (I) We identify Fodor and Pylyshyn’s main arguments in the context of the compositionality challenge for artificial neural networks and locate the nature of Lake and Baroni’s approach in refuting Fodor and Pylyshyn’s claims that neural networks cannot reliably develop compositional representations and structure-sensitive operations. (II) We show that within their setup, the model exhibits various non-systematic behaviors that can not be considered human-like and clearly violates structure-sensitivity. (III) We argue for several necessary aspects of training and evaluation of meta-learning systems to achieve and assess their systematicity, taking into account the relevance of compositional representations and structure-sensitive operations. (IV) We adapt the arguments of Fodor and Pylyshyn in light of the modern development of deep learning systems to argue for a future of models capable of learning symbolic representations for artificial cognition and representation learning.
@@ 3 Systematicity through Meta-Learning
[…]
@@ 3.1 Examining the lack of compositionality
Lake and Baroni mention that generalization beyond training occurs only with respect to new combinations of three grammar rules from the same set of grammar rules used during training. However, if we consider the invariance under the atomic assignments of colors to language tokens and the mere labeling of operations, we find that 179/200 validation episodes have a combination of non-primitive grammar operations that were already present in the 100000 training episodes. (See Appendix A.1 for details.) Thus, even if the model achieves highly systematic results on the test episodes, this could be due to memorization of the experienced operation patterns and learning to extract the correct labels from the episode’s support examples. However, we can even show that there is non-systematic behavior within their repository of test episodes; we re-evaluate their pre-trained model on the same set of test episodes, with the only difference that we did 10 evaluations of all query examples for each test episode, for statistical purposes, similar to the one episode they further evaluated against human performance. We find that the model performs worst on episodes #133, #32, and #122, with success rates of only 41%, 52%, and 54% on the query examples, respectively. (See the next paragraph and the Appendix A.2 for details.)
Failure in rule extraction. Further investigation of Episode #133 (see Table˜1) reveals that the model has trouble correctly processing the semantics of the language token with the hidden grammar rule and will therefore call it and confuse it up with the token (with ) which we will call . It seems to have a problem with the only example with , , which also happens to contain . But since has several iconic examples in the support, it is expected that a reasoner with compositional skills will be able to systematically use a single example and remain consistent with the rest of the support information. Considering the examples , , , human systematicity would at least suspect some semantics of that are different from those of .
@@
[…]
Violating structure-sensitivity. In addition to the previous two failure modes, which are related to the challenges of extracting information from the provided support examples, we also found query examples for episode #1 that reveal additional non-systematicity (see Table˜2 or Appendix A.2 for extended version). For queries with the patterns and we first see that the model never parses as intended. Instead of and , the stable output can be explained by parsing as intended. Instead of and – except for the cases, , , , where it would make no difference! Only the (also ambiguous) case is processed correctly in 6/10 cases – but with even worse performance than with the unambiguous examples. Despite the structural similarity to the other query examples, down to the color combination, we see a non-systematic deviation in the response, which raises doubts about compositional skills.
@@ 5 Alternative Views
6 Summary and Discussion
For this final section, we will reiterate the key points that make up our argument (see Sec. 1) and that we believe are important aspects of the goal of achieving meta-learning models capable of human-like systematic compositionality: (I) Criteria for compositionality. The main criteria for models with productive, systematic, and compositional capabilities remain compositional representation and structure-sensible operations. (II) Non-systematic behavior. Since Lake and Baroni’s model exhibits various non-systematic behaviors, it fails to demonstrate human-like compositional learning capabilities, and further refutes the presented claims that their meta-learning framework achieves human-like systematic generalization. (III) Assessment of compositionality. Systematic testing of multiple types of out-of-distribution episodes is necessary to assess compositional abilities and structure-sensitive operations. (IV) Emergence and Learning of Symbolic Representations. Meta-learning systems need to support the emergence of compositional symbolic representations during training. For this, we expect training tasks and model architectures that make iteration, self-validation, and self-correction over the extracted rule sets possible and necessary. The limitations of current neural models underscore the importance of hybrid architectures that integrate the strengths of symbolic and connectionist paradigms. Key advances in this direction include systematicity, reflective reasoning, and scalability.
@@ 7 Conclusion
While the importance of compositional representations and structure-sensitive operations for human-like systematicity remains, the previous consideration allows the training and testing of artificial neural networks that encourage the development of such properties. By bridging the gap between symbolic and connectionist principles, hybrid architectures may be particularly promising, since they do not suffer from the limitations of neural networks without symbolic machinery as specified by Fodor and Pylyshyn.
Overall, the continued relevance of Fodor and Pylyshyn’s critique underscores the challenges of developing systems capable of systematic generalization and compositional reasoning. While meta-learning frameworks represent significant progress, they do not address fundamental limitations. Future advances must embrace integrative approaches that combine the strengths of symbolic and connectionist paradigms, paving the way for a more robust understanding of artificial cognition. By addressing these challenges, we can move closer to realizing the vision of human-like artificial intelligence.
@@
This work was funded by the European Union (Grant Agreement no. 101120763 - TANGO). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Health and Digital Executive Agency (HaDEA). Neither the European Union nor the granting authority can be held responsible for them. The authors acknowledge the support of the German Science Foundation (DFG) research grant "Tractable Neuro-Causal Models" (KE 1686/8-1) as well as the AlephAlpha Collaboration lab 1141. This work was also supported by the Priority Program (SPP) 2422 in the subproject “Optimization of active surface design of high-speed progressive tools using machine and deep learning algorithms“ funded by the German Research Foundation (DFG). Further, we acknowledge support of the hessian.AI Service Center (funded by the Federal Ministry of Education and Research, BMBF, grant No 01IS22091), the “Third Wave of AI”, and “The Adaptive Mind”.
@@ A.1 Note on analysis for different combinations of non-primitive grammar operations
In section 3.1, we state that only 179/200 validation episodes have a combination of non-primitive grammar operations that were already present in the 100000 training episodes. This is the result of counting every different combination of 3 operations, unary () or binary (), present in both training and validation episodes, when abstracting the individual function names .
@@ A.2 Extended outputs for Lake and Baroni’s meta-learning testing episodes
Below, we include the full set of grammar rules, support examples, and query examples from Lake and Baroni’s meta-learning. We re-evaluated their pre-trained model on the same set of test episodes. Here we report the results for #133, #132, #122, and the modified #1.
@@
[…]
@@
[…]
@@
[…]
@@
[…]